OpenStudy (anonymous):

3(-2x-y)^2 equal to 3? F.x=-2 and y=1 G.x=-2 and y=-3 H.x=-1 and y=3 I.x=-1 and y=-3

OpenStudy (anonymous):

\[3\left( -2x-y \right)^2=3,\left( -2x-y \right)^2=\frac{ 3 }{ 3 }=1\] now check one by one.